# Quasi-stationary evolution of cubic-quintic NLSE drop-like solitons in DNA-protein systems

### v1

@@ 5 Conclusion
In this study, we investigate the nonlinear dynamics of DNA and the evolution of nonlinear excitations induced by protein interactions and the surrounding viscous medium, using a quasi-stationary approach. Unlike previous studies that model DNA as a rigid molecular chain governed by the NLSE, we employ a cubic-quintic NLSE to account for the flexibility of the DNA strands and the physiological properties of the protein-DNA binding system. We do not consider helicoidal terms in our model, as numerical simulations using the Yakushevich model of DNA have shown that their impact on soliton-like solutions is minimal. The viscous medium introduces damping effects on the nonlinear excitations, leading to a reduction in the amplitude and a constant velocity of the drop-like soliton. This damping becomes more pronounced as the damping coefficient $\epsilon$ increases,causing the soliton to decay more rapidly over time. In the case of pure internal DNA dynamics, where $\alpha_{1}=\alpha_{2}=0$, we can introduce various perturbations, such as periodic or pulse-like disturbances, depending on the nature of the double-stranded molecular system. These perturbations arise due to localized inhomogeneities in the DNA strands. In this work, we focus on the linear form of $R(\varphi)$, but similar analyses can be applied to other types of localized perturbations, not just in DNA but also in any physical system governed by the cubic-quintic NLSE. One of the key advantages of the quasi-stationary approach is its versatility in analyzing different soliton-like solutions, such as kinks, breathers, drops, and bubbles. While this approach provides a useful approximation for understanding the behavior of DNA’s energy and information encoded in the strands, a discrete analysis would be required for a more in-depth understanding of the underlying complexities of the double-stranded DNA system.

### v2

@@ 5 Conclusion
In this study, we analyzed the nonlinear dynamics of DNA and the evolution of nonlinear excitations induced by protein interactions and the surrounding viscous medium, using a quasi-stationary approach. Unlike previous studies that modeled DNA as a rigid molecular chain governed by the NLSE, we employed a cubic-quintic NLSE to account for the flexibility of DNA strands and the physiological properties of the protein-DNA binding system. We did not consider helicoidal terms in our model, as numerical simulations using the Yakushevich model have shown that their impact on soliton-like solutions is minimal. The viscous medium introduces damping effects on the nonlinear excitations, reducing the amplitude and resulting in a constant velocity for the drop-like soliton. As the damping coefficient $\epsilon$ increases, this damping effect becomes more pronounced, causing the soliton to decay more rapidly over time. For pure internal DNA dynamics, where $\alpha_{1}=\alpha_{2}=0$, we introduce various perturbations, such as periodic or pulse-like disturbances, which arise due to localized inhomogeneities in the DNA strands. While this work focuses on the linear form of $R(\varphi)$, similar analyses can be applied to other localized perturbations, both in DNA and in any system governed by the cubic-quintic NLSE. One of the key advantages of the quasi-stationary approach is its versatility in analyzing various soliton-like solutions, such as kinks, breathers, drops, and bubbles. Although this approach provides a useful approximation for understanding the behavior of DNA energy and information encoded in the strands, a more in-depth discrete analysis is necessary to fully capture the complexities of the double-stranded DNA system. Additionally, the direct impact of quasi-stationary evolution on compacton and anti-compacton solitons, which describe the open states in DNA chains, remains underexplored, presenting an opportunity for future research. Nevertheless, the present study provides explicit solutions for the direct analysis of these deformations.